# Linguistic Neuron Overlap Patterns to Facilitate Cross-lingual Transfer on Low-resource Languages

### v1

@@ 3.4 Bridge Language Selection
Based on the above observations, we quantify activation similarity between source-target overlap neurons and bridge-specific neurons and identify the optimal bridge language to facilitate X-ICL.
Given a language pair $L_{s}$ and $L_{t}$ and their overlap neurons $\mathcal{T}_{s,t}$, we obtain the activation value matrix $\mathbf{X}\in\mathbb{R}^{|\mathcal{T}_{s,t}|\times 2d}$ when probing LLMs with $d$ samples in both directions for balanced activation in $L_{s}$ and $L_{t}$. Simultaneously, we obtain $\mathbf{Y}\in\mathbb{R}^{|\bar{\mathcal{T}_{y}|}\times 2d}$, the activation value matrix for bridge language $L_{y}\in\mathcal{L}_{\text{candidate}}$, and $\bar{\mathcal{T}_{y}}=\mathcal{T}_{y}-\mathcal{T}_{s,t}-\mathcal{T}_{y^{\prime}}$ represents language-specific neurons in $L_{y}$ excludes neurons in $\mathcal{T}_{s,t}$ and $\mathcal{T}_{y^{\prime}}$, where $L_{y^{\prime}}\neq L_{y}$.
We employ the Hilbert-Schmidt Independence Criterion (HSIC) Gretton et al. (2005) to measure the nonlinear dependency between activation matrices of $\mathbf{X}$ and $\mathbf{Y}$. Average pooling will be performed to standardize matrices of $\mathbf{X}$ and $\mathbf{Y}$ to have the same row dimension $n$. The formal HSIC is calculated as: $\text{HSIC}(\mathbf{X},\mathbf{Y})=n^{-2}\text{Tr}(\mathbf{KHLH})$, where $\mathrm{Tr}(\cdot)$ is the trace operation, $\mathbf{K},\mathbf{L}\in\mathbb{R}^{n\times n}$ are learned kernel matrices for $\mathbf{X}$ and $\mathbf{Y}$. $\mathbf{H}=\mathbf{I}_{n\times n}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$ is a centering matrix, where $\mathbf{I}_{n\times n}$ is the identity matrix of size $n\times n$, $\mathbf{1}_{n}$ is a vector of $n$ ones. Rather than computing HSIC over the entire activation matrices, we adopt a bidirectional maximum matching strategy to measure the strongest dependency between individual neurons in one set and the entire distribution of the other, computed as:
$\mathrm{HSIC}(\bar{\mathcal{T}}_{y},\mathcal{T}_{s,t})=$ (3)
$\frac{1}{2}\left(\max_{i}\mathrm{HSIC}(\mathbf{x}_{i},\mathbf{Y})+\max_{j}\mathrm{HSIC}(\mathbf{X},\mathbf{y}_{j})\right)$
where $\mathbf{x}_{i}\in\mathbf{X},\mathbf{y}_{j}\in\mathbf{Y}$. We compute the dependency scores per layer, then average them across middle $K$ layers to get $L_{y}$’s selection probability:
$p(L_{y}|L_{s}\text{-}L_{t})=\frac{1}{K}\sum_{i=1}^{K}\text{HSIC}\left(\bar{\mathcal{T}}_{y}(i),\mathcal{T}_{s,t}(i)\right)$ (4)
where $K$ is determined according to embedding semantic similarity and discussed in section B.3. Finally, the optimal bridge $L^{*}$ is selected by:
$L^{*}=\arg\max_{L_{y}\in\mathcal{L}_{\text{candidate}}}p(L_{y}|L_{s}\text{-}L_{t})$ (5)
@@ 4.1 Experiment Setup
Implementation. We evaluate BridgeX-ICL on $2$ cross-lingual tasks and $15$ languages covering $7$ diverse language families: Indo-European: English (En), German (De), French (Fr), Italian (It), Portuguese (Pt), Spanish (Es); Uralic: Finnish (Fi), Hungarian (Hu); Afro-Asiatic: Arabic (Ar), Hebrew (He); Austronesian: Indonesian (Id), Tagalog (Tl); Sino-Tibetan: Chinese (Zh); Japonic: Japanese (Ja); Niger-Congo: Swahili (Sw).
@@
The evaluation focuses on LLMs’ cross-lingual transfer on low-resource languages. Therefore, we take He, Tl, Sw, and Ja as target languages to build $15$ cross-lingual pairs, covering moderate-to-low (e.g., Ar-Sw), and high-to-low (e.g., En-He) pairs, both within and across families. The classification of high-, moderate-, and low-resource languages is based on their proportion in LLM’s training corpora, following previous work Cieri et al. (2016). Since the bridge language should be well supported by LLMs, we take 6 languages in Indo-European family as candidate bridges.
Datasets. To evaluate the generalization of bridge selection beyond the BLI task, we also evaluate on an additional cross-lingual task: Machine Reading Comprehension (MRC) using the Belebele dataset Bandarkar et al. (2024).
@@
To evaluate low-resource languages, a key challenge lies in lacking evaluation benchmarks. Although the ground truth MUSE Conneau et al. (2017) provides $110$ bilingual dictionaries for BLI task, it does not cover the tested 15 language pairs. To solve this, we used English as a pivot to build $L_{s}$-$L_{t}$ dictionary from $L_{s}$-English and $L_{t}$-English. For languages not in MUSE (e.g., Swahili), we extracted word pairs from wiktionary_bli Izbicki (2022) to build En-Sw. We checked all word pairs using both Google and Microsoft translators to ensure quality and selected $1,000$ word pairs for each language pair that are consistently validated by both systems. The constructed BLI dictionaries are available at: https://anonymous.4open.science/r/BLI--0481/.
@@
Metrics. For BLI task, we use Precision@N metric, which measures the accuracy of the model’s top $N$ candidate translations. In this study, $N$ is set to 1. For the MRC task, we use accuracy to measure whether the model selects the correct answer from multiple choices.
@@
LLMs. We conducted experiments on two open-source LLMs: LLaMA-3-8B Grattafiori et al. (2024) and Mistral-7B-Instruct-v0.3 Jiang et al. (2023). Their training corpora cover 176 and 53 languages, respectively, which include all the experimental low-resource languages and allow us to explore the underlying linguistic mechanisms.
[…]
@@ 4.2.1 LLMs’ Linguistic Spectrum Discussion
(a) LLaMA 3’s Linguistic spectrum
(b) Human language similarity from Glottolog
Figure 4: Comparison of linguistic spectrum calculated based on overlap neurons in LLaMA 3 and human language similarity derived from Glottolog Phylogenetic Trees, including 15 languages from 7 families. Darker blue indicates a higher language similarity.
This section discusses the linguistic similarities across $15$ languages from $7$ families, calculated based on overlap neurons in LLaMA 3 and Mistral, as presented in Figure 4 (a) and Figure 8 in B.2, respectively. To evaluate how closely the linguistic spectrum learned by LLMs align with that of human languages, we leverage Glottolog Phylogenetic Trees Hammarström et al. (2023), which encode hierarchical relationships among 8,000+ human languages, to draw human language similarity in Figure 4(b). The detailed linguistic similarity computed based on Glottolog is in Appendix A. In Figure 4, color intensity represents degree of similarity between languages, with darker blue indicating stronger degree and the diagonal is self-similarity (1.0).
Table 1: Comparison of BLI task improvement on 15 language pairs. The highest gains are marked with bold in few-shot and zero-shot with bridge methods. ’-’ indicates the selected bridge is either the source or target language.
@@
Human source +2.80 +6.70 +3.80 +1.50 -9.50 - -7.40 -0.60 -11.90 -3.20 - -3.90 +12.30 -3.30 -1.70
Human target +2.80 +9.60 -0.10 +5.00 -0.50 - -7.10 - -12.10 +3.80 - +1.20 +16.40 -2.30 +2.30
@@
Human source -3.60 -0.20 +0.90 +0.90 -8.70 - -2.20 +0.80 -5.70 -5.70 - -0.40 -0.10 +2.30 +2.10
Human target -3.60 -1.60 -0.60 +0.20 -9.50 - +4.90 - -4.70 -7.20 - -0.20 -9.00 +1.40 +1.20
@@
Linguistic spectrum learned by LLMs are not fully aligned with human languages. We can observe a strong neural similarity within language families, marked using red text box in Figure 4, which matches human linguistic taxonomy. For example, within Afro-Asiatic family, both from Arabic (Ar) and Hebrew (He), have a high neuron overlap (0.470), greater than Arabic-Swahili with 0.161. In addition, high-resource Indo-European languages, such as Fr-It and Pt-Es, show the highest overlap scores, with darkest blue in the bottom-right corner of the heatmap. But this alignment breaks down between high- and low-resource languages, like Arabic-French with similarity $0.479$.
LLMs build their own distinct understanding of language relationships. The calculated linguistic spectrum of LLaMA 3 and Mistral are similar but not the same. The two models may choose different bridges for the same language pair, as discussed later. This is likely because their internal linguistic relationships are primarily shaped by the distribution of languages in training corpora, as noted in Philippy et al. (2023). This explains why Arabic has the strongest similarity ($0.479$) with French in Romance family, rather than with Hebrew ($0.470$) in the same Afro-Asiatic.
@@ 4.2.2 Cross-lingual Results Analysis
This section compares performance of BridgeX-ICL against various baselines on BLI task across $15$ language pairs, as presented in Table 1.
Table 2: Evaluation on MRC cross-lingual task. Red color highlights the different bridge selection and bold marks the highest gains at each language pair.
@@
We can find: 1) LLMs exhibit poor and imbalanced performance on low-resource languages. For example, LLaMA 3 achieves its best BLI performance of $69.90$ on Ar-Ja pair, but worst of $25.90$ on Id-Sw pair. 2) LLMs are few-shot multilingual learners. However, few-shot X-ICL does not consistently yield stable gains, e.g., one-shot falling below zero-shot, and will not lead to further gains when the number of shots goes beyond $3$. It indicates when applying few-shot X-ICL, 3-shots will be enough. 3) Zero-shot with bridge is a simple yet data-efficient strategy for low-resource languages. BridgeX-ICL finds $9$ optimal bridges out of $15$ language pairs, achieving average performance of two-shot X-ICL across all pairs, followed by the English-bridge method. While human source/target methods are the least effective. It seems using English as the default bridge is cost-effective, which will be discussed in section 5.1.
5 Discussion
5.1 Application of Bridge Language
Beyond BLI task, Table 2 evaluates BridgeX-ICL on the MRC cross-lingual task. The prompt for zero-shot with bridge in MRC is detailed in Appendix E. Results show our approach still works well on MRC task and benefits more on LLaMA 3 than Mistral. For example, BridgeX-ICL improves performance of LLaMA 3 by an average of 6.036.03% over the zero-shot baseline across 15 language pairs, while the average improvement is 4.484.48% on Mistral.
English is selected as the optimal bridge in 9 out of 15 language pairs. The reason is partly due to LLMs’ unbalanced language abilities across 5 candidate bridges. As discussed in Figure 3 another key factor is LLMs’s inherent preference of English-pivot during cross-lingual transfer.

### v2

@@ 3.4 Bridge Language Selection
Based on the above observations, we leverage the constructed probe data with $d$ samples per language pair to identify the optimal bridge language $L_{*}$ to facilitate X-ICL from source language $L_{s}$ to target language $L_{t}$.
Given $L_{s}$ and $L_{t}$ and their overlap neurons $\mathcal{T}_{s,t}$ identified in section 3.3.2, we obtain the activation value matrix $\mathbf{X}\in\mathbb{R}^{|\mathcal{T}_{s,t}|\times 2d}$ by prompting the LLM with $d$ samples in both directions for balanced neuron activation in $L_{s}$ and $L_{t}$. We also obtain the activation matrix $\mathbf{Y}\in\mathbb{R}^{|\bar{\mathcal{T}_{y}|}\times 2d}$ for a candidate bridge language $L_{y}\in\mathcal{L}_{\text{candidate}}$. Here, $\bar{\mathcal{T}_{y}}=\mathcal{T}_{y}-\mathcal{T}_{s,t}-\mathcal{T}_{y^{\prime}}$ represents the set of language-specific neurons in $L_{y}$, excluding those shared with $\mathcal{T}_{s,t}$ or with any other language $L_{y^{\prime}}\neq L_{y}$.
We employ HSIC Gretton et al. (2005) to measure the nonlinear dependency between activation matrices $\mathbf{X}$ and $\mathbf{Y}$. Average pooling will be performed to standardize matrices of $\mathbf{X}$ and $\mathbf{Y}$ to have the same row dimension $n$. The HSIC is formally calculated as: $\text{HSIC}(\mathbf{X},\mathbf{Y})=n^{-2}\text{Tr}(\mathbf{KHLH})$, where $\mathrm{Tr}(\cdot)$ is the trace operation, $\mathbf{K},\mathbf{L}\in\mathbb{R}^{n\times n}$ are learned kernel matrices for $\mathbf{X}$ and $\mathbf{Y}$. $\mathbf{H}=\mathbf{I}_{n\times n}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$ is a centering matrix, where $\mathbf{I}_{n\times n}$ is the identity matrix of size $n\times n$, $\mathbf{1}_{n}$ is a vector of $n$ ones. Rather than computing HSIC over the entire activation matrices, we adopt a bidirectional maximum matching strategy that measures the strongest dependency between a single neuron vector $\mathbf{x}_{i}\in\mathbf{X}$ ($\mathbf{y}_{j}\in\mathbf{Y}$) and the entire distribution of the other, where $\mathbf{x}_{i},\mathbf{y}_{i}\in\mathbb{R}^{d}$, computed as:
$\mathrm{H}(\bar{\mathcal{T}}_{y},\mathcal{T}_{s,t})=\frac{1}{2}\Big(\max_{i}\,\mathrm{HSIC}(\mathbf{x}_{i},\mathbf{Y})+\max_{j}\,\mathrm{HSIC}(\mathbf{X},\mathbf{y}_{j})\Big)$ (3)
We compute the dependency scores layer by layer and average them across the middle $K$ layers to estimate the selection probability of $L_{y}$:
$p(L_{y}|L_{s}\rightarrow L_{t})=\frac{1}{K}\sum_{i=1}^{K}\text{H}\left(\bar{\mathcal{T}}_{y}(i),\mathcal{T}_{s,t}(i)\right)$ (4)
where $K$ is determined according to embedding semantic similarity and discussed in section C.3. Finally, the optimal bridge $L^{*}$ is selected by:
$L_{*}=\arg\max_{L_{y}\in\mathcal{L}_{\text{candidate}}}p(L_{y}|L_{s}\rightarrow L_{t})$ (5)
@@ 4.1 Experiment Setup
Implementation. We evaluate BridgeX-ICL on $4$ cross-lingual tasks and $15$ languages covering $7$ diverse language families: Indo-European: English (En), German (De), French (Fr), Italian (It), Portuguese (Pt), Spanish (Es); Uralic: Finnish (Fi), Hungarian (Hu); Afro-Asiatic: Arabic (Ar), Hebrew (He); Austronesian: Indonesian (Id), Tagalog (Tl); Sino-Tibetan: Chinese (Zh); Japonic: Japanese (Ja); Niger-Congo: Swahili (Sw).
@@
As the evaluation focuses on LLMs’ cross-lingual transfer on low-resource languages, we take He, Tl, Sw, and Ja as target languages to build $15$ cross-lingual pairs, covering moderate-to-low (e.g., Ar-Sw) and high-to-low (e.g., En-He) pairs, both within and across language families. The classification of high-, moderate-, and low-resource languages is based on their proportion in LLMs’ training corpora, following previous work Cieri et al. (2016). Since the bridge language should be well supported by LLMs, we select $6$ languages in the Indo-European family as candidate bridges. We also conducted an exploratory experiment using bridges not in the Indo-European family discussed in section 5.1.
Datasets. To evaluate the generalization of bridge selection beyond the BLI task, we further evaluate cross-lingual Machine Reading Comprehension (MRC) using the Belebele dataset Bandarkar et al. (2024). To verify robustness, we additionally consider two cross-lingual tasks: Cross-Lingual Question Answering (CLQA) and Cross-Lingual Natural Language Inference (XNLI). Since CLQA and XNLI cover only six of our evaluated language pairs, their experimental results are reported in Appendix B.
@@
To evaluate low-resource languages, a key challenge lies in the lack of evaluation benchmarks. Although the ground-truth MUSE Conneau et al. (2017) provides $110$ bilingual dictionaries for the BLI task, it does not cover the $15$ language pairs we tested. To address this, we used English as a pivot to build $L_{s}$-$L_{t}$ dictionary from $L_{s}$-English and $L_{t}$-English. For languages not in MUSE (e.g., Swahili), we extracted word pairs from wiktionary_bli Izbicki (2022) to build En-Sw. We verified all word pairs using both Google and Microsoft translators to ensure quality and selected $1,000$ word pairs for each language pair that are consistently validated by both systems  [^1] The constructed BLI dictionaries are available at: https://github.com/xuyuemei/BLI-..
@@
Metrics. For the BLI task, we use the Precision@N metric, which measures the accuracy of the model’s top-$N$ candidate translations. In this study, $N$ is set to $1$. For the MRC task, we use accuracy to evaluate whether the model selects the correct answer from multiple choices.
@@
LLMs. We conducted experiments on two open-source LLMs: LLaMA-3-8B Dubey et al. (2024) and Mistral-7B-Instruct-v0.3 Jiang et al. (2023). Their training corpora cover 176 and 53 languages, respectively, which include all the experimental low-resource languages and allow us to explore the underlying linguistic mechanisms.
Baselines. Baselines are divided into zero-shot, few-shot, and zero-shot with bridge. Specifically, zero-shot is the basic prompt setup, and few-shot builds on the zero-shot prompt by adding 1, 2, 3, or 4 samples. For zero-shot with bridge approach, we compare BridgeX-ICL with 5 baselines described below. 1) Phylogenetic Distance Source/Target (Ph.D Source or Ph.D Target): Select the bridge language closest to the source or target language according to language genealogy of Glottolog Trees Hammarström et al. (2023). 2) English Bridge: Use English as the bridge language. 3) Sharing Matters: Wang et al.(2024) used activation values to find shared neurons across languages. We select language with the most shared neurons as the bridge language. 4) IoU: Use Intersection over Union (IoU) Tan et al. (2024), also known as Jaccard index, to measure linguistic distance. Given neuron sets $\mathcal{T}_{u},\mathcal{T}_{v}$ associated with language $L_{u}$, $L_{v}$, $\mathrm{IoU}(\mathcal{T}_{u},\mathcal{T}_{v})=|\mathcal{T}_{u}\cap\mathcal{T}_{v}|/|\mathcal{T}_{u}\cup\mathcal{T}_{v}|$. Language with the highest average IoU score to $L_{s}$ and $L_{t}$ is selected. 5) LAPE_overlap: Use entropy-based LAPE Tang et al. (2024) to identify language-specific neurons. We then compute cosine similarity on overlap neurons between the bridge and source/target. The language with the highest average similarity is selected.
(a) Linguistic spectrum in LLaMA 3
(b) Language similarity from Glottolog Trees
Figure 4: Comparison of linguistic spectrum calculated based on overlapping neurons in LLaMA 3 and language similarity derived from Glottolog Phylogenetic Trees, including $15$ languages from 77 families. Darker blue indicates a higher language similarity.
@@ 4.2.1 LLMs’ Linguistic Spectrum Discussion
This section discusses the linguistic similarities across $15$ languages from $7$ families, calculated based on overlapping neurons in LLaMA 3 and Mistral, as presented in Figure 4(a) and Figure 8 in Appendix C.2, respectively. To evaluate how closely the linguistic spectrum learned by LLMs aligns with that of human languages, we leverage Glottolog Phylogenetic Trees Hammarström et al. (2023), which encode hierarchical relationships among 8,000+ human languages, to derive human language similarity in Figure 4(b). The detailed process for computing linguistic similarity based on Glottolog Trees is in Appendix A. In Figure 4, darker blue indicates stronger similarity, and the diagonal denotes self-similarity (1.0).
Linguistic spectrum learned by LLMs is not fully aligned with human language phylogeny. We observe strong neural similarities within language families, highlighted by a red text box in Figure 4, which matches human linguistic taxonomy. For example, Arabic (Ar) and Hebrew (He), within the Afro-Asiatic family, exhibit a high neuron overlap (0.470), greater than Arabic-Swahili with 0.161 similarity. A similar pattern appears with Indonesian (Id) and Tagalog (Tl), both from the Austronesian family. In addition, high-resource Indo-European languages, such as French-Italian (Fr-It) and Portuguese-Spanish (Pt-Es), show the highest overlap scores, with the darkest blue in the bottom-right corner of the heatmap. But this alignment breaks down in high-to-low resource language pairs, and some cross-family pairs display unexpected high similarity scores, likely reflecting training data distribution rather than intrinsic linguistic relationships.
LLMs build their own distinct understanding of language relationships. The calculated linguistic spectra of LLaMA 3 and Mistral are similar but not identical. The two models may choose different bridges for the same language pair, as discussed later. We observe that Arabic has the strongest similarity (0.479) with French in the Romance family, rather than with Hebrew (0.470) from its own Afro-Asiatic family. This counterintuitive result is likely due to the linguistic relationships learned by LLMs being primarily shaped by the distribution of languages in training corpora, as noted in Philippy et al. (2023).
Table 1: Comparison of BLI task improvement on $15$ language pairs. The highest gains are marked with bold in few-shot and zero-shot with bridge methods. ‘-’ indicates the selected bridge is either the source or target language.
@@
Ph.D Source +2.80 +6.70 +3.80 +1.50 -9.50 - -7.40 -0.60 -11.90 -3.20 - -3.90 +12.30 -3.30 -1.70
Ph.D Target +2.80 +9.60 -0.10 +5.00 -0.50 - -7.10 - -12.10 +3.80 - +1.20 +16.40 -2.30 +2.30
@@
Ph.D Source -3.60 -0.20 +0.90 +0.90 -8.70 - -2.20 +0.80 -5.70 -5.70 - -0.40 -0.10 +2.30 +2.10
Ph.D Target -3.60 -1.60 -0.60 +0.20 -9.50 - +4.90 - -4.70 -7.20 - -0.20 -9.00 +1.40 +1.20
@@ 4.2.2 Cross-lingual Results Analysis
Table 1 compares the performance of BridgeX-ICL against various baselines on the BLI task across $15$ language pairs.
We observe the following findings: 1) LLMs exhibit poor and imbalanced performance on low-resource languages. For example, LLaMA 3 achieves its best BLI performance of $69.90$ on the Ar-Ja pair, but its worst of $25.90$ on the Id-Sw pair. 2) LLMs are few-shot multilingual learners, as also noted in Al Nazi et al. (2024). However, few-shot X-ICL does not consistently yield stable gains. For example, one-shot sometimes performs worse than zero-shot, and performance often stabilizes or may decline once the number of shots exceeds $3$. This suggests that when applying few-shot X-ICL, 3-shots will be enough. 3) Zero-shot with bridge is a simple yet data-efficient strategy for low-resource languages. BridgeX-ICL finds $9$ optimal bridges out of $15$ language pairs, achieving average performance of two-shot X-ICL across all pairs, followed by the English-bridge method. While phylogenetic distance source/target methods are the least effective. It seems using English as the default bridge is cost-effective, which will be discussed in section 5.2.
5 Discussion
5.1 Candidate Bridge Language Selection
To validate the rationale for selecting Indo-European languages as bridge candidates, Table 2 presents an exploratory experiment in which all languages in our study were evaluated as potential bridges, using $6$ representative language pairs.
It shows that Indo-European languages on average outperform the nine non-Indo-European languages. Interestingly, some non-Latin-script languages, such as Chinese, also demonstrate potential as effective bridges. The results provided valuable insights: only languages well supported by LLMs functioned as effective bridges. Based on this observation, we selected $6$ languages in Indo-European as the final candidates in our study.
Table 2: Candidate bridge language selection comparing Indo-European languages with 99 non-Indo-European languages. ’-’ indicates the selected bridge is either the source or target language.
Zero-shot Zh Ja Ar He Id Tl Fi Hu Sw Indo-European (Avg.)
Zh-Ja 67.10 - - +3.30 +0.80 0 -5.10 +2.80 +5.20 -9.90 +9.43
Zh-He 44.10 - +7.90 +9.60 - +4.80 -0.20 +6.70 +10.80 -8.80 +14.37
Ar-Tl 46.70 +4.50 +3.10 - -7.40 -7.10 - +3.40 +4.90 -6.10 +6.88
Ar-He 47.00 +14.90 +11.70 - - -5.20 +1.40 +11.60 +15.30 -3.00 +15.72
Id-Ja 62.50 -2.40 - -9.00 -5.20 - -11.90 -12.10 -9.60 -24.80 +1.40
Id-Sw 25.90 +3.20 +1.50 +1.20 -0.80 - -3.90 -0.50 +1.30 - +3.43
5.2 Application of Bridge Language
Beyond the BLI task, Table 3 evaluates BridgeX-ICL on the MRC cross-lingual task. The prompt for zero-shot with bridge in MRC is detailed in Appendix F. Appendix B further evaluates BridgeX-ICL on CLQA and XNLI cross-lingual tasks. Results show our approach works well and benefits more from LLaMA 3 than from Mistral. For example, BridgeX-ICL improves the performance of LLaMA 3 by an average of 6.036.03% over the zero-shot baseline across $15$ language pairs, while the average improvement on Mistral is 4.484.48%.
English is selected as the optimal bridge in $9$ out of $15$ language pairs in LLaMA 3 ($6$ out of $15$ in Mistral). This is partly due to the unbalanced language abilities of LLMs across 5 candidate Indo-European bridges. As discussed in Figure 3, another key factor is the model’s inherent preference for English-pivot during cross-lingual transfer.
Table 3: Evaluation on MRC cross-lingual task. Red color highlights the different bridge selections, and bold marks the highest gains at each language pair.